A random variable Z has the following cumulative distribution function

1 for x>=10

.6 for 6<=x<10

.4 for 4<=x<6

.3 for 1<=x<4

.2 for 1<=x<3

0 for x<1

Use the cumulative distribution function directly to find

A. P(3<=x<=6)

B. P(3<x<6)
Answer #1

A. P(3 x 6) = P(x 6) - P(x < 3)

= 0.6 - 0.2

= 0.4

B. P(3 < x < 6) = P(x < 6) - P(x < 3)

= 0.4 - 0.2

= 0.2